TETSXPREME:

help i forgot how 2 do this xd

TETSXPREME:

1 attachment
Hulu:

:)

Hulu:

times the whole by the bottom number and add the top to the numberator of the new number.

Hulu:

so our first number would be 65/12

Hulu:

and the second number would be 62/11

Hulu:

compare.

KjSaif:

If you look at it 7/11 is closer to a whole then 5/12.

TETSXPREME:

so first one 65?

KjSaif:

that being said would give you the idea that \[5 \frac{ 7 }{ 11 }\] is greater then \[5 \frac{ 5 }{ 12 }\]

KjSaif:

tet imagine a pie chart.

KjSaif:

with 12 slices in one pie and 11 in another pie.

TETSXPREME:

12x11?= 132?

KjSaif:

uh not quite lol. lets go back to the looking at the fraction

JamesTDG:

just convert the fractions into decimals and you are good to go

KjSaif:

ok look tet

1 attachment
KjSaif:

this is 7/11

KjSaif:

this is 5/12

1 attachment
KjSaif:

which one is closer to being completely shaded

TETSXPREME:

close to what?

KjSaif:

grey, shaded .__.

KjSaif:

see the pics i sent ?

TETSXPREME:

yea

KjSaif:

you see how the 7/11 circle is closer to being completely grey ?

KjSaif:

@kjsaif wrote:
that being said would give you the idea that \[5 \frac{ 7 }{ 11 }\] is greater then \[5 \frac{ 5 }{ 12 }\]

TETSXPREME:

thank you soo much

Hulu:

Np c;

KjSaif:

np B)